OpenStudy (anonymous):

Find each point of discontinuity of f , and for each give the value of the point of discontinuity and evaluate the indicated one-sided limits. Let f(x)=x^2+9/9-x^2

myininaya (myininaya):

f is discontinuous at x=3 and x=-3 x->3^(-), then f(x)->+inf x->3^(+), then f(x)->-inf x->-3^(-), then f(x)->-inf x->-3^(+), then f(x)->+inf

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

hey myininaya do you have a sec to think about something?

myininaya (myininaya):

maybe its not too hard right my belly hurts

OpenStudy (anonymous):

points discontinuity: 9-x^2=0 which gives x=3 & x=-3

myininaya (myininaya):

what do you want me to think about

OpenStudy (anonymous):

i don't know if it's hard or not but I can't see how to do it right now. I want to expand n!/(n-k)!

OpenStudy (anonymous):

write it as a series possibly

myininaya (myininaya):

ok let me think and stuff doesn't sound too hard but it might be

OpenStudy (anonymous):

i figured it would've been done already but couldn't find an expansion anywhere

myininaya (myininaya):

i think i should assume n>k

OpenStudy (anonymous):

yeah we can assume that

OpenStudy (anonymous):

it's the term in front of x^n when you take the kth derivative: (n)(n-1)(n-2)(n-3)...

OpenStudy (anonymous):

n(n-1)(n-2)...(n-k+1)x^(n-k)

myininaya (myininaya):

if we have \[\frac{n!}{(n-8)!}= \frac{n!}{(n-0)!(n-1)(n-2)(n-3)(n-4)....(n-8)} =\frac{1}{(n-1)(n-2)(n-3)(n-4)....(n-8)}\] so if we have \[\frac{n!}{(n-k)!} =\frac{n!}{(n-(k-k))!(n-(k-(k+1)))...(n-(k-3))(n-(k-2))(n-(k-1))(n-k)}\] =\[\frac{1}{(n-(k-(k+1)))...(n-(k-3))(n-(k-2))(n-(k-1))(n-k)}\]{}\]

myininaya (myininaya):

does this help?

myininaya (myininaya):

my thingy got cutoff lol

OpenStudy (anonymous):

(n-8)!=(n-8)(n-9)(n-10)...

myininaya (myininaya):

yeah it got cut off

OpenStudy (anonymous):

where did the (n)(n-1)(n-2).. terms come from

myininaya (myininaya):

oh wait no i didn't include those terms

OpenStudy (anonymous):

i need a representation of the polynomial created by multiplying terms together: (n)(n-1)(n-2)=(n^2-n)(n-2)=n^3-3n^2+2n

OpenStudy (anonymous):

so the first term always has coeff 1. the last term is (-1)^(k+1)(k-1)! the second term is k(k-1)/2

OpenStudy (anonymous):

so i need a general formula for whichever term of the polynomial i need

OpenStudy (anonymous):

like what is the coefficient on n^3 when k =6?

myininaya (myininaya):

\[\frac{5!}{(5-2)!}=\frac{5!}{5!(5-1)(5-2)}=\frac{1}{(5-1)(5-2)}=\frac{1}{4*3}=\frac{1}{12}\] \[\frac{5!}{3!}=\frac{5*4*3*2*1}{3*2*1}=5*4=20\] not equal im dumb :(

OpenStudy (anonymous):

i think its a harder problem than it sounds at first I cant figure it out

myininaya (myininaya):

1/12 not equal to 5!/(5-2)! contradiction im my logic which makes it non logic which makes it retarded

OpenStudy (anonymous):

haha its just a mistake

myininaya (myininaya):

which makes me a retard

OpenStudy (anonymous):

ive made some bad mistakes in front of a class before

myininaya (myininaya):

ok sometimes it helps me to work on non general cases before i go to a general case

OpenStudy (anonymous):

i hate making mistakes too though

OpenStudy (anonymous):

k try looking at n*(n-1)(n-2)(n-3) if you want and see if you can find a nice form for the coefficients maybe

OpenStudy (anonymous):

i know when all the terms are the same we can use combinatorics to find the coefficients

OpenStudy (anonymous):

so maybe its some extension of that

myininaya (myininaya):

so we want to write a summation right not one those weird big pi symbols that mean multiply

OpenStudy (anonymous):

summation where each part is a term of the polynomial

OpenStudy (anonymous):

i think it will help me find a general solution to this: \[\sum_{n=0}^{\infty}\frac{n^k}{e^{an}}\]

OpenStudy (anonymous):

i know how to do it for specific k and a but I want a general solution. I think there will be a somewhat nice one

myininaya (myininaya):

did you try putting it in wolfram i don't think this will be nice looking

OpenStudy (anonymous):

with k and a as variables or for a specific value of k and a?

OpenStudy (anonymous):

i tried to get maple to solve it generally and it wouldnt do it

OpenStudy (anonymous):

the specific solutions are nice though

OpenStudy (anonymous):

i dont know how to type it into wolfram

myininaya (myininaya):

http://www.wolframalpha.com/input/?i=n!%2F(n-k)!

OpenStudy (anonymous):

what is a hurwitz lurch transcendent!?

OpenStudy (anonymous):

thats the solution apparently haha

myininaya (myininaya):

lol i have no clue never heard of it

OpenStudy (anonymous):

hmm maybe there isnt a nice solution to this

OpenStudy (anonymous):

well depending on what a hurwitz lurch transcendent is i geuss

myininaya (myininaya):

lol that series was for n=0 it wasn't even a general case for all n but for k

myininaya (myininaya):

i can't look at this anymore lol

OpenStudy (anonymous):

haha im reading about hurwitz function on wikipedia it's horrible.

OpenStudy (anonymous):

allright maybe i will try a different problem

myininaya (myininaya):

good luck

OpenStudy (anonymous):

thanks for looking at it. im off to bed talk to you later :)

myininaya (myininaya):

i go to bed too :) goodnight